OpenStudy (anonymous):

What is the y intercept of y= -3/2x + 4

jimthompson5910 (jim_thompson5910):

Similar Example: y = 2/3x - 7 is in the form y = mx + b because you can write y = 2/3x - 7 as y = 2/3x + (-7) So you can see that b = -7 The value of b is the y-intercept, so because b = -7, the y-intercept is -7 which is the point (0, -7) Again, this is just an example and NOT the answer to your specific question

OpenStudy (anonymous):

So the Y intercept for y= -3/2x + 4 would be 4?

jimthompson5910 (jim_thompson5910):

you got it

jimthompson5910 (jim_thompson5910):

which is the point (0, 4)

OpenStudy (anonymous):

Thanks again! :D

jimthompson5910 (jim_thompson5910):

sure thing